OpenStudy (anonymous):

Solve for x. -3 + 8x - 5 = -8

OpenStudy (blurbendy):

Start by getting the x variable alone on one side. do you know how to do this?

OpenStudy (anonymous):

Not really

OpenStudy (blurbendy):

Okay, since 8x is already on the left side, let's keep it that way. So, we need to get everything that doesn't have an x variable to the right side.

OpenStudy (blurbendy):

we can start with the -3 -3 + 8x -5 = -8 (Add three to both sides) 8x - 5 = -5 Do you see what I did?

OpenStudy (anonymous):

Moved them around

OpenStudy (blurbendy):

pretty much, all I did was add 3 to both sides. this cancelled the 3 on the left, and -8 + 3 = -5, which is how I ended up with -5 on the right

OpenStudy (blurbendy):

Can you tell me what to do next with 8x -5 = -5 ?

OpenStudy (anonymous):

I'm at a 5th grade level for math..sooo. no

OpenStudy (blurbendy):

okay, we're going to do the same thing. but we're going to add 5 to both sides to get the x variable by itself 8x - 5 = -5 8x = 0 (because -5 + 5 = 0) make sense?

OpenStudy (anonymous):

Not at all, sorry I have a really hard time understanding math

OpenStudy (blurbendy):

it's okay. what don't you understand?

OpenStudy (anonymous):

....Everything

OpenStudy (blurbendy):

well the equation we started with was -3 + 8x - 5 = -8 and we want to know what 'x' is equal to. so we have to solve for it. that's all we're doing but to do that, we need to get the 'x' by itself

OpenStudy (blurbendy):

let me write the equation we started with differently.

OpenStudy (blurbendy):

8x -3 -5 = -8 all i did was move the -3 next to the -5

OpenStudy (blurbendy):

now, what is -3 -5 ?

OpenStudy (anonymous):

-8?

OpenStudy (blurbendy):

perfect. so our equation becomes 8x -8 = -8

OpenStudy (blurbendy):

now the 'x' needs to be by itself, so what is the opposite of -8?

OpenStudy (anonymous):

8

OpenStudy (blurbendy):

yes, so we are going to ADD 8 to both sides 8x -8 = -8 +8 +8

OpenStudy (blurbendy):

what is -8 + 8 ?

OpenStudy (anonymous):

16

OpenStudy (blurbendy):

-8 + 8 equals 0, not 16 it's the same as 8 - 8 which is 0 do you see?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

The cancel each other out

OpenStudy (blurbendy):

yes, good job. so, we end up with 8x = 0

OpenStudy (blurbendy):

almost done, we need the x by itself, but there is 8 in front of it. so what do we do? we divide both sides by 8

OpenStudy (blurbendy):

what is 8/8 ?

OpenStudy (anonymous):

1or 0

OpenStudy (blurbendy):

it's 1. ANY number divided by itself is ALWAYS 1

OpenStudy (blurbendy):

and what is 0 / 8 ?

OpenStudy (anonymous):

0?

OpenStudy (blurbendy):

perfect. so we end up with 1x = 0 which you can write as just x = 0

OpenStudy (anonymous):

Thank you

OpenStudy (blurbendy):

you can check that your answer is right by putting it back into the equation 8(0) -3 -5 = -8 -8 = -8 which is true.

OpenStudy (blurbendy):

you did good, good job!

OpenStudy (anonymous):

Lol thanks